OpenStudy (anonymous):

Needing actual help with converting units of measure , not just give me the answer and move on. ++Medal & Fan -I have example questions -Thanks.

OpenStudy (anonymous):

The distance between New York City, New York, and Los Angeles, California, is a approx 4485 kilometers. What is the distance between these two major cities in miles? --Im asking you to help me set up proportions.

OpenStudy (bibby):

in dimensional analysis, you want to set things up so that you can "cancel" the units. like so: 320 km to miles: \(\huge \frac{320\cancel{km}}{}*\frac{1mi}{1.60934\cancel{km}}=whatever\)

OpenStudy (anonymous):

sooo \[\frac{4485 kilometers}{} \times \frac{ 1 mile }{ 1.60934 kilometers } = 4795.9899?\]

OpenStudy (bibby):

you did something weird, multiplying by a fraction shouldn't make something bigger

OpenStudy (anonymous):

Bibby is correct; do it like this: 4485km * 1mile = 4485km Then Divide: 4485/1.60934miles which gives you 2,786 miles approximately. You can check my multiplying the above number by 1.60934, which gives you 4485 km! Enjoy!

OpenStudy (anonymous):

I don't know, they're teaching me to write out the number given in the problem (in this case 4485km) the multiply that by whatever err how to i explain this. Like here's an example: 5 pints = ___ liters? 5 pints * 1 quart/2 pints * 1 liter/1.057 quarts = 5/8 or 0.625

OpenStudy (bibby):

the last line is what we just did, is that not what they want?

OpenStudy (anonymous):

OHHHHH I GOT IT duh. I'm sorry. I was multiplying the numerator by the denominator instead of multiplying across and dividing the fraction

OpenStudy (anonymous):

Yes, it is what we did. Just show your units. When doing these questions, try to eliminate variables you don't want and leave the variable you want to solve. Notice for your question we crossed out km, but left miles.

OpenStudy (bibby):

:D

OpenStudy (anonymous):

Thanks guys ^^

OpenStudy (anonymous):

Sure thing!